Hi. I continually have difficulty solving distance problems involving multiple variables. Here is my question. Thank you so much.
A railroad and a highway intersect at right angles. A train is 10 miles from the intersection at the same moment that a car is 6 miles from the intersection. Both are traveling at 30 mph. How long until they are 4 miles apart?
There are two answers:
First time = _______ minutes
Second time = ______ minutes

Answer by stanbon(75887)

A railroad and a highway intersect at right angles. A train is 10 miles from the intersection at the same moment that a car is 6 miles from the intersection. Both are traveling at 30 mph. How long until they are 4 miles apart?
----
Draw the picture so you can see what is happening with the distances:
Train DATA:
rate = 30 mph ; time = t hours ; distance = 30t miles
---
Car DATA:
rate = 30 mph ; time = t hour ; distance = 30t miles
-----
Distance train is from the intersection: 10-30t miles
Distance car is from the intersection: 6-30t miles
-----
Use Pythagoras:
(10-30t)^2 + (6-30t)^2 = 4^2
100 - 600t + 900t^2 + 36 - 360t + 900t^2 = 16
-------
1800t^2 -960t + 120 = 0
-------
Divide thru by 120 to get:
15t^2 - 8t + 1 = 0
(3t-1)(5t-1) = 0
----
t = 1/3 hr = 20 min
Or
t = 1/5 hr = 12 min
====================
Cheers,
Stan H.
====================